OpenStudy (chris215):

The slope of the tangent to a curve at any point (x, y) on the curve is x/y . Find the equation of the curve if the point (2, -3) is on the curve.

OpenStudy (chris215):

I got x^2 + y^2 = 13

OpenStudy (princeharryyy):

let equation of tangent => y = mx +c given slope = x/y at point (x,y) so eq => y = (x/y)*x +c point ((2, -3)) is on this line insert (2, -3) in place of x,y find c

OpenStudy (chris215):

x^2-y^2=a^2 2^2+3^2=a^2 13=a^2 x^2 + y^2 = 13

OpenStudy (princeharryyy):

were you able to find small c

OpenStudy (princeharryyy):

hold on. your c will be -5/3 and equation will be y^2 = x^2 -3*5/(-3) = y^2 = x^2 +5

OpenStudy (chris215):

ohhh ok so my answer would be x^2-y^2=5

OpenStudy (princeharryyy):

y^2-x^2 = 5

OpenStudy (chris215):

thank you!!